How do you work this?
the sum of two and three times the difference of a number and five is equal to four times the differnce a number and one

Answer by mananth(16949)

let number be x
3(x-5)+2 =4(x-1)
3x-15+2=4x-4
3x-13=4x-4
3x-4x=13-4
-x=9
divide by -1
x=-9
m.ananth@hotmail.ca